OpenStudy (anonymous):

Solve the equation or formula for the indicated variable. s=5r^2t, for t A t = s/5 - r B t = 25r/s C t = r^2-5s D t =s/5r^2

jimthompson5910 (jim_thompson5910):

Similar Problem: \[\Large s = 10p^2r\] Solve for 'r', you do this by dividing both sides by 10p^2 \[\Large s = 10p^2r\] \[\Large \frac{s}{10p^2} = r\] \[\Large r = \frac{s}{10p^2}\]

OpenStudy (anonymous):

thank you :) D correct?

jimthompson5910 (jim_thompson5910):

you got it